how do you solve this equation 9x² - 4x - 5 = 0
Answer by checkley77(12844)

9x² - 4x - 5 = 0
(9x+5)(x-1)=0
9x+5=0
9x=-5
x=-5/9 ans.
x-1=0
x=1 ans.